# Languages of Boundedly-Ambiguous Vector Addition Systems with States

### v1

@@ 3 Tools for separating BAmb and NonDet
In this section we develop two techniques for showing that a language is not recognized by a $k$-ambiguous VASS: Lemma 3.2 and Lemma 3.23.
@@ 3.1 Dominating block
The first tool is based on the following observation about a language not recognized by a $k$-ambiguous VASS.
@@
For every $k\in\mathbb{N}_{+}$ language
$L_{k}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid\exists_{1\leq i\leq k+1}n_{i}\geq n_{i+1}\}$
is not recognized by a $k$-ambiguous VASS.
@@
Let $\Sigma$ be an alphabet such that $b\notin\Sigma$ and let $L$ be a language over $\Sigma$. For each function $f:L\rightarrow\mathbb{N}_{\omega}$ such that $\sup{f}=\omega$ (recall that $\sup{f}=\sup{\{f(x)\mid x\in L\}}$) language $L_{1}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}bw\mid w\in L,\exists_{1\leq i\leq k+1}n_{i}\geq n_{i+1}\lor n_{k+2}\geq f(w)\}$ is not recognized by a $k$-ambiguous VASS.
@@
Let us fix $k\in\mathbb{N}_{+}$ and let $L=\{\varepsilon\}$. Let $f:L\rightarrow\mathbb{N}_{\omega}$ be defined as $f(\varepsilon)=\omega$. Hence by Lemma 3.2 we get that $L_{k}$ is not recognized by a $k$-ambiguous VASS.
Below we sketch the idea behind the proof of Lemma 3.2, most of the proof is delegated to the Appendix. We assume, for the sake of contradiction, that $L_{1}$ is recognized by a $k$-ambiguous VASS and aim at a contradiction by showing $k+1$ different runs over the same word. We first consider $k+1$ words $w_{1},\ldots,w_{k+1}\in L_{1}$, where $u\in L$ is a particularly chosen word and $N_{0}<N_{1}<\ldots<N_{k+2}\in\mathbb{N}$ are particularly chosen constants:
@@
Then we dive into combinatorics of VASS runs $\rho_{i}$ over words $w_{i}$ and conclude that there are specific pumping cycles in $\rho_{i}$. We formulate below one of the used lemmas in order to illustrate the kind of arguments we consider and, as it is also used in Section 3.2.
@@
Let $L$ be a language over $\Sigma=\{a,b\}$ recognized by some $k$-ambiguous $d$-VASS $V$. For each $n\in\mathbb{N}$ there exists a constant $C$ such that each run $\rho$ in $V$ such that:
@@
    For each $j\in[n]$ let $A_{j}=\bigcup_{1\leq i<j}\textup{support}(\beta_{i})$ then $\beta_{j}$ is nonnegative on counters from $[d]\setminus A_{j}$
@@
    For each $j\in[n]$ there is no $\lambda$, $\delta$ and $\lambda^{\prime}$ such that $\delta$ is a nonnegative loop on counters from $[d]\setminus A_{j}$, $\lambda^{\prime}\neq\varepsilon$ and $\lambda\delta\lambda^{\prime}=\alpha_{j}\beta_{j}$
@@
and $n_{j}=2m^{k+3-j}\Pi_{l=j}^{k+2}N_{l}!$ (where $m$ is the maximal norm of a transition). More concretely, $\rho^{\prime}_{i}$ result from pumping the loops $\beta_{i}$ in $\rho_{i}$ more times. A challenge it to show that the resulting $\rho^{\prime}_{i}$ are indeed different, which we achieve by more careful, but technical investigation of the runs. Existence of $k+1$ different runs over the same word $w$ is a contradiction with the assumption that $L_{1}$ is recognized by a $k$-ambiguous VASS and finishes the proof.
@@
Our plan for the proof of Lemma 3.2 is to assume, for the sake of contradiction, that $L_{1}$ is recognized by a $k$-ambiguous VASS and reach a contradiction by showing $k+1$ different runs over the same word. For this we will use two pumping techniques: Lemma 3.5 and Lemma 3.4. Hence, we present first a lemma, which gives conditions for finding a specific loop in a run. Observe, that this lemma is general and is not restricted to $k$-ambiguous VASSs.
Lemma 3.5.
@@
Formally, we extend firing a transition from configurations from the set $Q\times\mathbb{N}^{d}$ to configurations from the set $Q\times\mathbb{N}_{\omega}^{d}$ by setting, that for each $k\in\mathbb{Z}$ we have $\omega+k=\omega$. Now let us observe the following:
@@
Now we are ready to prove Lemma 3.5.
Proof 3.7 (Proof of Lemma 3.5).
For each state $q$ of $V$ we construct the domination tree from configuration $q(u)$ where $u=(u_{1},u_{2},\ldots,u_{d})$ where for $i\in S$ we have $u_{i}=n$ and for $i\notin S$ we have $u_{i}=\omega$. As constant $M$ we take the maximal depth of such trees plus one. Observe, that we do not have to check all options $u_{i}=k\leq n$ (because a sequence of transitions which can be fired from some configuration can also be fired from a bigger configuration) and $M$ is well-defined because of the Claim 1. Assume, towards contradiction, that there exists a run $\rho$ of length at least $M$, starting from a configuration $q_{1}(v)$ in which values of counters from $S$ are at most $n$ in which there is no loop, which has a non-negative effect on counters from $S$. Because for $i\in[d]$ we have $v_{i}\leq u_{i}$ the same sequence of transitions can be fired from $q_{1}(u)$. Let $T$ be the domination tree constructed from $q_{1}(u)$. We can follow run $\rho$ in $T$ upon reaching a leaf of the tree. Let $\lambda$ be the prefix of $\rho$, which reached the leaf in the tree. Observe, that $|\lambda|<M$. By the construction of the domination tree we know, that there are two possibilities:
@@
Beside Lemma 3.5 we also need another tool for characterization of the structure of runs of $k$-ambiguous VASS, namely Lemma 3.4, which was formulated in Section 3. Here we present its proof.
Proof 3.8 (Proof of Lemma 3.4).
We prove this lemma by induction on $n$. For $n=0$ we have $\rho=\varepsilon$ and hence all decomposition conditions are satisfied. Now we assume, that this Lemma is true for $n-1$ and we show that this implies Lemma 3.4 for $n$.
@@
Recall the definition of set $A_{n}$ from condition 5, let $m$ be the maximal norm of a transition, $s$ be the maximal norm of an initial configuration and let us define constants $T=s+2(n-1)mC$ and $C_{n}=M(V,A_{n},T)$ given by Lemma 3.5. Let us also define constant $K_{n}=\max_{S\subseteq[d]}M(V,S,T+mC_{n})$. Let also $C^{\prime}=\max(C+1,C_{n},K_{n})$. This will be a constant for Lemma 3.4 and $n$. We have two cases. The first case is that $\pi_{2}=\alpha_{n}\beta_{n}^{a_{n}}\alpha_{n}^{\prime}$ for some loop $\beta_{n}$, which is nonnegative on counters from $[d]\setminus A_{n}$ and $|\alpha_{n}\beta_{n}|\leq C_{n}$ and $\beta_{n}$ is not a prefix of $\alpha_{n}^{\prime}$. The second case is that such a decomposition of $\pi_{2}$ does not exist. We start the proof with the second case. Case 2. Observe, that for each $i\in[n-1]$ and $l\in[d]\setminus A_{n}\subseteq[d]\setminus A_{i}$:
@@
satisfies all the conditions. We start with conditions 2-6 as Condition 1 is the most challenging to prove. Condition 2. For $j<n-1$ this condition follows from the induction assumption. For $j=n-1$ we have that $\alpha_{n-1}^{\prime}=\lambda t$. Recall that $t$ was a transition reading letter $b$ and from the induction assumption $\lambda$ does not read letter $b$. Moreover, $\alpha_{n}^{\prime}$ is a suffix of $\pi_{2}$, which does not read the letter $b$. Hence the condition follows. Condition 3. For $j<n$ the condition follows from the induction assumption and for $j=n$ we have that $\alpha_{n}\beta_{n}$ is a prefix of $\pi_{2}$, which does not read the letter $b$. Condition 4. For $j<n$ we have this condition from induction assumption and for $j=n$ we have that $\beta_{n}\neq\varepsilon$ is a loop. Condition 5. This condition for $j<n$ follows from inductions assumption and for $j=n$ follows from the fact that $\beta_{n}$ is a nonnegative loop on counters from $[d]\setminus A_{n}$ as required. Condition 6. For $j<n$ we have it from induction assumption. For $j=n$ we have it from minimality of $|\alpha_{n}\beta_{n}|$. Condition 1. Observe, that for $i\in[n-2]$ from induction assumption we have $|\alpha_{i}|,|\beta_{i}|,|\alpha_{i}^{\prime}|\leq C\leq C^{\prime}$ and moreover $|\alpha_{n-1}|,|\beta_{n-1}|\leq C\leq C^{\prime}$. Moreover, from induction assumption, $|\alpha_{n-1}|,|\beta_{n-1}|\leq C\leq C^{\prime}$ and $|\alpha_{n-1}^{\prime}|=|\lambda t|\leq C+1\leq C^{\prime}$. Additionally $|\alpha_{n}\beta_{n}|\leq C_{n}\leq C^{\prime}$. Therefore it is enough to prove that $|\alpha_{n}^{\prime}|\leq K_{n}\leq C^{\prime}$. Let
@@
Recall that $K_{n}=M(V,S,T+mC_{n})$ and $T=2(n-1)mC$, so if $|\alpha_{n}^{\prime}|>K_{n}$ then $\alpha_{n}^{\prime}$ contains a nonnegative loop on counters from $S$. Hence it is enough to prove, that $\alpha_{n}^{\prime}$ does not contain such loop. Assume, for the sake of contradiction, that $\alpha_{n}^{\prime}=\lambda\gamma\lambda^{\prime}$ such that $\gamma$ is a nonnegative loop on counters from $S$. We define accepting runs $\phi_{1},\phi_{2},\ldots,\phi_{k+1}$, which will read the same word, but will be different and hence this will create a contradiction with the fact that $V$ is $k$-ambiguous VASS. Intuitively, due to loops $\beta_{n}$ and $\gamma$ we have more degrees of freedom in the last block and we can create $k+1$ different runs over the same word. As $\rho$ is a prefix of an accepting run we know, that there exists $\rho_{1}$ such that $\rho\rho_{1}$ is an accepting run. Firstly, we define runs $\psi_{1}$, $\psi_{2}$ and $\psi_{3}$, which will be parts of $\phi_{1},\phi_{2},\ldots,\phi_{k+1}$:
@@
Proof 3.12 (Proof of Lemma 3.2).
The proof is by contradiction. We assume that $L_{1}$ is recognized by a $k$-ambiguous VASS $V$. Firstly we define a few constants. Then we define word $w_{1},w_{2},\ldots,w_{k+1}\in L_{1}$ and respective accepting runs $\rho_{1},\rho_{2},\ldots,\rho_{k+1}$. Then we decompose each of these runs using Claim 6 and apply pumping to each of the runs. Finally we get $k+1$ different runs $\rho_{1}^{\prime},\rho_{2}^{\prime},\ldots,\rho_{k+1}^{\prime}$ over the same word, which contradicts the fact that $V$ is $k$-ambiguous .
Assume, for the sake of contradiction, that $L_{1}$ is recognized by a $k$-ambiguous VASS $V$ and let $m$ be the maximal norm of a transition in $V$ (w.l.o.g. $m\geq 2$). Let also $n$ be the maximal norm of an initial configuration. Let us fix a few constants. Let $N_{0}$ be equal to $12C+$ 1 (where $C$ is the constant given by Lemma 3.4). Moreover, for $i\in[k+2]$ we define $N_{i}$ as $\max(N_{i-1}+2,M(V,[d],M_{i}))$ where $M_{i}=\Sigma_{j=1}^{i-1}m(N_{j}!+1)+n$ and $M(V,[d],M_{i})$ is the constant given by Lemma 3.5 for VASS $V$, subset of counters $[d]$, and initial value of the counters $M_{i}$. Observe, that because of this for each $i\in[k+1]$ we have $N_{i+1}!\geq(N_{i}+2)!>(N_{i}+1)!$ and for each $i\in[k+2]$ we have $N_{i}\geq 3$ (because $N_{0}\in\mathbb{N}_{+}$). Let also $u$ be a word from $L$ such that $f(u)\geq 2N_{k+2}!$. Such word exists because $\sup{f}=\omega$.
@@
Observe, that $w_{i}\in L$ because $N_{i}!\geq N_{0}!$ and let $\rho_{i}$ be an accepting run over $w_{i}$. The intuition is, that $\rho_{i}$ has a nonnegative loop on all the counters in each part reading a block of letters a possibly except the one reading $i+1$ block. This is formalized in the following claim:
@@
Now we have a nonnegative loop in each block possibly except the $i+1$ block. We want to find a loop in this block, which we will be able to pump, probably with the help of previous nonnegative loops in the earlier blocks. For this, we need Lemma 3.4. We formalize our goal in the following claim, which is similar to Lemma 3.4, but more specific to our situation. In particular conditions 1, 2, 3 and 4-5 from Claim 6 correspond respectively to conditions 2, 3, 4 and 5 from Lemma 3.4.
@@
    We have that $\beta_{i+1}^{i}$ is a nonnegative loop on counters from $[d]\setminus\bigcup_{1\leq j\leq i}\textup{support}(\beta_{j}^{i})$, $|\beta_{i+1}^{i}|\leq N_{0}$ and $\beta_{i+1}^{i}$ is not a nonnegative loop on all the counters
Proof 3.14.
@@
Observe, that for simplicity we drop upper index $i$. Clearly conditions 1 and 2 are satisfied. Now we show, that the other conditions are also satisfied. We start with conditions 3 and 5, which are significantly simpler to prove than Condition 4. Condition 3. Because of Lemma 3.4 we have to only show that $\beta_{j}^{i}\neq\varepsilon$. Observe, that the number of letter a in the $j$-th block equals at least $N_{0}!\geq N_{0}=C$ (where $C$ is constant from Lemma 3.4). Hence $\beta_{j}^{i}\neq\varepsilon$. Condition 5. Because of Lemma 3.4 we have that $\beta_{i+1}$ is a nonnegative loop on counters from $[d]\setminus\bigcup_{1\leq j\leq i}\textup{support}(\beta_{j})$. Moreover, we also have:
@@
Proof 3.15.
Recall that $m$ is the maximal norm of a transition and set $b=b_{j}m|\beta_{j}^{i}|$ and let us fix $l\in[d]$. If $\textsc{eff}_{l}(\beta_{j})\geq 0$ the inequality from the Claim 6 holds. Hence we can assume $\textsc{eff}_{l}(\beta_{j})<0$. Because $\beta_{j}$ is a nonnegative loop on the counters from $[d]\setminus\bigcup_{1\leq m<j}\textup{support}(\beta_{m}^{i})$ there exists $n<j$ such that $\textsc{eff}_{l}(\beta_{n})>0$. Hence
@@
Therefore we reached a contradiction with the fact that $V$ is a $k$-ambiguous VASS and hence $L_{1}$ is not recognized by a $k$-ambiguous VASS, which finishes our proof.

### v2

@@ 3 Tools for separating BAmb and NonDet
In this section we develop two techniques for showing that a language is not recognised by a $k$-ambiguous VASS: Lemma 3.2 and Lemma 3.20.
@@ 3.1 Dominating block
The first tool is based on the following observation about a language not recognised by a $k$-ambiguous VASS.
@@
For every $k\in\mathbb{N}_{+}$ the language
$L_{k}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid\exists_{i\in[k+1]}\ n_{i}\geq n_{i+1}\}$
is not recognised by a $k$-ambiguous VASS.
@@
Let $\Sigma$ be an alphabet such that $b\notin\Sigma$ and let $L$ be a language over $\Sigma$. For each function $f:L\rightarrow\mathbb{N}_{\omega}$ such that $\sup{f}=\omega$ (recall that $\sup{f}=\sup{\{f(x)\mid x\in L\}}$) language $L_{1}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}bw\mid w\in L,\exists_{i\in[k+1]}\ n_{i}\geq n_{i+1}\lor n_{k+2}\geq f(w)\}$ is not recognised by a $k$-ambiguous VASS.
@@
Let us fix $k\in\mathbb{N}_{+}$ and let $L=\{\varepsilon\}$. Let $f:L\rightarrow\mathbb{N}_{\omega}$ be defined as $f(\varepsilon)=\omega$. Hence by Lemma 3.2 we get that $L_{k}$ is not recognised by a $k$-ambiguous VASS.
Below we sketch the idea behind the proof of Lemma 3.2. The whole proof can be found in the Appendix. We assume, for the sake of contradiction, that $L_{1}$ is recognised by a $k$-ambiguous VASS and aim at a contradiction by showing $k+1$ different runs over the same word. We first consider $k+1$ words $w_{1},\ldots,w_{k+1}\in L_{1}$, where $u\in L$ is a particularly chosen word and $N_{0}<N_{1}<\ldots<N_{k+2}\in\mathbb{N}$ are particularly chosen constants:
@@
Then we dive into combinatorics of VASS runs $\rho_{i}$ over words $w_{i}$ and conclude that there are specific pumping cycles in $\rho_{i}$. We show two main lemmas, which analyse the structure of the runs. The first one gives conditions for finding a specific loop in a run. Observe, that this lemma is general and is not restricted to $k$-ambiguous VASSs.
@@
The proof of Lemma 3.4, which can be found in the Appendix, uses rather standard techniques similar to the ones in the Karp-Miller construction solving the coverability problem for VASS [coverability_tree]. Lemma 3.4 is used to find loops in words $w_{i}$. For example one can show that if $N_{1}$ is big enough then one can find a loop as in Lemma 3.4 in the first block of letters $a$. Similarly, if $N_{2}$ is big enough wrt. to $N_{1}$, one can find an appropriate loop in the second block of letters $a$. Using this approach one can find loops in blocks of letters $a$, if the block is much longer than the whole prefix before it. This is however not true for the block of length $N_{0}$. Therefore we need a more sophisticated tool in order to be able to modify runs over all the words $w_{i}$ to runs over the same word.
This requires the more challenging part of the proof, which provides a detailed characterisation of runs of $k$-ambiguous VASS. We formulate below one of the used lemmas in order to illustrate the kind of arguments we consider and, as it is also used in Section 3.5.
Lemma 3.5.
Let $L$ be a language over $\Sigma=\{a,b\}$ recognised by some $k$-ambiguous $d$-VASS $V$. For each $n\in\mathbb{N}$ there exists a constant $C$ such that each run $\rho$ in $V$ such that:
@@
    For each $j\in[n]$ let $A_{j}=\bigcup_{1\leq i<j}\textup{support}(\beta_{i})$ then $\beta_{j}$ is nonnegative on counters from $d]\setminus A_{j}$
@@
    For each $j\in[n]$ there is no $\lambda$, $\delta$ and $\lambda^{\prime}$ such that $\delta$ is a nonnegative loop on counters from $d]\setminus A_{j}$, $\lambda^{\prime}\neq\varepsilon$ and $\lambda\delta\lambda^{\prime}=\alpha_{j}\beta_{j}$
The proof of Lemma 3.5 uses the Lemma 3.4 and a lot of combinatorial observations. It is moved to the Appendix.
@@
and $n_{j}=2m^{k+3-j}\Pi_{l=j}^{k+2}N_{l}!$ (where $m$ is the maximal norm of a transition). More concretely, $\rho^{\prime}_{i}$ is obtained by pumping the loops $\beta_{i}$ in $\rho_{i}$ more times. A challenge it to show that the resulting $\rho^{\prime}_{i}$ are indeed different, which we achieve by more careful, but technical investigation of the runs. Existence of $k+1$ different runs over the same word $w$ is a contradiction with the assumption that $L_{1}$ is recognised by a $k$-ambiguous VASS and finishes the proof.
@@
3.2 Proof of Lemma 3.4
@@
Formally, we extend firing a transition from configurations from the set $Q\times\mathbb{N}^{d}$ to configurations from the set $Q\times\mathbb{N}_{\omega}^{d}$ by setting, that for each $k\in\mathbb{Z}$ we have $\omega+k=\omega$. Now let us observe the following claim, which can be proven by standard techniques using well quasi orders and Kőnig’s Lemma [konig].
@@
Now we are ready to prove Lemma 3.4. For each state $q$ of $V$ we construct the domination tree from configuration $q(u)$ where $u=(u_{1},u_{2},\ldots,u_{d})$ where for $i\in S$ we have $u_{i}=n$ and for $i\notin S$ we have $u_{i}=\omega$. As constant $M$ we take the maximal depth of such trees plus one. Observe, that we do not have to check all options $u_{i}=k\leq n$ (because a sequence of transitions which can be fired from some configuration can also be fired from a bigger configuration) and $M$ is well-defined because of the Claim 1. Assume, towards contradiction, that there exists a run $\rho$ of length at least $M$, starting from a configuration $q_{1}(v)$ in which values of counters from $S$ are at most $n$ in which there is no loop, which has a non-negative effect on counters from $S$. Because for $i\in[d]$ we have $v_{i}\leq u_{i}$ the same sequence of transitions can be fired from $q_{1}(u)$. Let $T$ be the domination tree constructed from $q_{1}(u)$. We can follow run $\rho$ in $T$ upon reaching a leaf of the tree. Let $\lambda$ be the prefix of $\rho$, which reached the leaf in the tree. Observe, that $|\lambda|<M$. By the construction of the domination tree we know, that there are two possibilities:
@@
3.3 Proof of Lemma 3.5
We prove this lemma by induction on $n$. For $n=0$ we have $\rho=\varepsilon$ and hence all decomposition conditions are satisfied. Now we assume, that this Lemma is true for n1n-1 and we show that this implies Lemma 3.5 for $n$.
@@
Recall the definition of set $A_{n}$ from condition 5, let $m$ be the maximal norm of a transition, $s$ be the maximal norm of an initial configuration and let us define constants T=s+2(n1)mCT=s+2(n-1)mC and $C_{n}=M(V,A_{n},T)$ given by Lemma 3.4. Let us also define constant $K_{n}=\max_{S\subseteq[d]}M(V,S,T+mC_{n})$. Let also $C^{\prime}=\max(C+1,C_{n},K_{n})$. This will be a constant for Lemma 3.5 and $n$. We have two cases. The first case is that $\pi_{2}=\alpha_{n}\beta_{n}^{a_{n}}\alpha_{n}^{\prime}$ for some loop $\beta_{n}$, which is nonnegative on counters from $d]\setminus A_{n}$ and $|\alpha_{n}\beta_{n}|\leq C_{n}$ and $\beta_{n}$ is not a prefix of $\alpha_{n}^{\prime}$. The second case is that such a decomposition of $\pi_{2}$ does not exist. We start the proof with the second case. Case 2. Observe, that for each $i\in[n-1]$ and $l\in[d]\setminus A_{n}\subseteq[d]\setminus A_{i}$:
@@
satisfies all the conditions. We start with conditions 2-6 as Condition 1 is the most challenging to prove. Condition 2. For j<n1j<n-1 this condition follows from the induction assumption. For j=n1j=n-1 we have that $\alpha_{n-1}^{\prime}=\lambda t$. Recall that $t$ was a transition reading letter $b$ and from the induction assumption $\lambda$ does not read letter $b$. Moreover, $\alpha_{n}^{\prime}$ is a suffix of $\pi_{2}$, which does not read the letter $b$. Hence the condition follows. Condition 3. For $j<n$ the condition follows from the induction assumption and for $j=n$ we have that $\alpha_{n}\beta_{n}$ is a prefix of $\pi_{2}$, which does not read the letter $b$. Condition 4. For $j<n$ we have this condition from induction assumption and for $j=n$ we have that $\beta_{n}\neq\varepsilon$ is a loop. Condition 5. This condition for $j<n$ follows from inductions assumption and for $j=n$ follows from the fact that $\beta_{n}$ is a nonnegative loop on counters from $d]\setminus A_{n}$ as required. Condition 6. For $j<n$ we have it from induction assumption. For $j=n$ we have it from minimality of $|\alpha_{n}\beta_{n}|$. Condition 1. Observe, that for $i\in[n-2]$ from induction assumption we have $|\alpha_{i}|,|\beta_{i}|,|\alpha_{i}^{\prime}|\leq C\leq C^{\prime}$ and moreover $|\alpha_{n-1}|,|\beta_{n-1}|\leq C\leq C^{\prime}$. Moreover, from induction assumption, $|\alpha_{n-1}|,|\beta_{n-1}|\leq C\leq C^{\prime}$ and $|\alpha_{n-1}^{\prime}|=|\lambda t|\leq C+1\leq C^{\prime}$. Additionally $|\alpha_{n}\beta_{n}|\leq C_{n}\leq C^{\prime}$. Therefore it is enough to prove that $|\alpha_{n}^{\prime}|\leq K_{n}\leq C^{\prime}$. Let
@@
Recall that $K_{n}=M(V,S,T+mC_{n})$ and T=2(n1)mCT=2(n-1)mC, so if $|\alpha_{n}^{\prime}|>K_{n}$ then $\alpha_{n}^{\prime}$ contains a nonnegative loop on counters from $S$. Hence it is enough to prove, that $\alpha_{n}^{\prime}$ does not contain such loop. Assume, for the sake of contradiction, that $\alpha_{n}^{\prime}=\lambda\gamma\lambda^{\prime}$ such that $\gamma$ is a nonnegative loop on counters from $S$. We define accepting runs $\phi_{1},\phi_{2},\ldots,\phi_{k+1}$, which will read the same word, but will be different and hence this will create a contradiction with the fact that $V$ is $k$-ambiguous VASS. Intuitively, due to loops $\beta_{n}$ and $\gamma$ we have more degrees of freedom in the last block and we can create $k+1$ different runs over the same word. As $\rho$ is a prefix of an accepting run we know, that there exists $\rho_{1}$ such that $\rho\rho_{1}$ is an accepting run. Firstly, we define runs $\psi_{1}$, $\psi_{2}$ and $\psi_{3}$, which will be parts of $\phi_{1},\phi_{2},\ldots,\phi_{k+1}$:
@@
Since we have proven Claims 23 and 4 we know that runs $\phi_{i}$ are all accepting, different and read the same word. This concludes the proof of Lemma 3.5.
@@
3.4 Proof of Lemma 3.2
The proof is by contradiction. We assume that $L_{1}$ is recognised by a $k$-ambiguous VASS $V$. Firstly we define a few constants. Then we define word $w_{1},w_{2},\ldots,w_{k+1}\in L_{1}$ and respective accepting runs $\rho_{1},\rho_{2},\ldots,\rho_{k+1}$. Then we decompose each of these runs using Claim 6 and apply pumping to each of the runs. Finally we get $k+1$ different runs $\rho_{1}^{\prime},\rho_{2}^{\prime},\ldots,\rho_{k+1}^{\prime}$ over the same word, which contradicts the fact that $V$ is $k$-ambiguous .
Assume, for the sake of contradiction, that $L_{1}$ is recognised by a $k$-ambiguous VASS $V$ and let $m$ be the maximal norm of a transition in $V$ (w.l.o.g. $m\geq 2$). Let also $n$ be the maximal norm of an initial configuration. Let us fix a few constants. Let $N_{0}$ be equal to $2C+1$ (where $C$ is the constant given by Lemma 3.5). Moreover, for $i\in[k+2]$ we define $N_{i}$ as $\max(N_{i-1}+2,M(V,[d],M_{i}))$ where $M_{i}=\Sigma_{j=1}^{i-1}m(N_{j}!+1)+n$ and $M(V,[d],M_{i})$ is the constant given by Lemma 3.4 for VASS $V$, subset of counters $[d]$, and initial value of the counters $M_{i}$. Observe, that because of this for each $i\in[k+1]$ we have $N_{i+1}!\geq(N_{i}+2)!>(N_{i}+1)!$ and for each $i\in[k+2]$ we have $N_{i}\geq 3$ (because $N_{0}\in\mathbb{N}_{+}$). Let also $u$ be a word from $L$ such that $f(u)\geq 2N_{k+2}!$. Such word exists because $\sup{f}=\omega$.
@@
Observe, that $w_{i}\in L$ because $N_{i}!\geq N_{0}!$ and let $\rho_{i}$ be an accepting run over $w_{i}$. The intuition is, that $\rho_{i}$ has a nonnegative loop on all the counters in each part reading a block of letters a possibly except the one reading $i+1$ block. This is formalised in the following claim:
@@
Now we have a nonnegative loop in each block possibly except the $i+1$ block. We want to find a loop in this block, which we will be able to pump, probably with the help of previous nonnegative loops in the earlier blocks. For this, we need Lemma 3.5. We formalise our goal in the following claim, which is similar to Lemma 3.5, but more specific to our situation. In particular conditions 1, 2, 3 and 4-5 from Claim 6 correspond respectively to conditions 2, 3, 4 and 5 from Lemma 3.5.
@@
    We have that $\beta_{i+1}^{i}$ is a nonnegative loop on counters from $d]\setminus\bigcup_{1\leq j\leq i}\textup{support}(\beta_{j}^{i})$, $|\beta_{i+1}^{i}|\leq N_{0}$ and $\beta_{i+1}^{i}$ is not a nonnegative loop on all the counters
Proof 3.11.
@@
Observe, that for simplicity we drop upper index $i$. Clearly conditions 1 and 2 are satisfied. Now we show, that the other conditions are also satisfied. We start with conditions 3 and 5, which are significantly simpler to prove than Condition 4. Condition 3. Because of Lemma 3.5 we have to only show that $\beta_{j}^{i}\neq\varepsilon$. Observe, that the number of letter a in the $j$-th block equals at least $N_{0}!\geq N_{0}=C$ (where $C$ is constant from Lemma 3.5). Hence $\beta_{j}^{i}\neq\varepsilon$. Condition 5. Because of Lemma 3.5 we have that $\beta_{i+1}$ is a nonnegative loop on counters from $d]\setminus\bigcup_{1\leq j\leq i}\textup{support}(\beta_{j})$. Moreover, we also have:
@@
Proof 3.12.
Recall that $m$ is the maximal norm of a transition and set $b=b_{j}m|\beta_{j}^{i}|$ and let us fix $l\in[d]$. If $\textsc{eff}_{l}(\beta_{j})\geq 0$ the inequality from the Claim 6 holds. Hence we can assume $\textsc{eff}_{l}(\beta_{j})<0$. Because $\beta_{j}$ is a nonnegative loop on the counters from $d]\setminus\bigcup_{1\leq m<j}\textup{support}(\beta_{m}^{i})$ there exists $n<j$ such that $\textsc{eff}_{l}(\beta_{n})>0$. Hence
@@
Therefore we reached a contradiction with the fact that $V$ is a $k$-ambiguous VASS and hence $L_{1}$ is not recognised by a $k$-ambiguous VASS, which finishes the proof of Lemma 3.2.